OpenStudy (anonymous):

trouble with this problem also "Catherine is camping along a river. it takes her 1.5 hours to paddle her canoe 6 miles upstream from her campsite. Catherine turns her canoe around and returns 6 miles downstream to her campsite in exactly 1 hour. what is the rate of the rivers current and the rate of Catherine's paddling in still water?"

OpenStudy (lgbasallote):

All you need to know is: downstream = speed of boat + speed of current upstream = speed of boat - speed of current Therefore... let x = speed of boat y = speed of current DOWNSTREAM: 6mph = x + y UPSTREAM: 4 mph = x -y Systems of Linear Equations.... x + y = 6 x - y = 4 2x = 10 x = 5 The speed of the boat is 5 mph...The speed of the current is 1 mph.

OpenStudy (mertsj):